OpenStudy (anonymous):

How do you solve 4 to power of x plus 2= five to power of x

OpenStudy (anonymous):

\[4^{x+2}=5^{x}\]

OpenStudy (anonymous):

is that your question

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\ln 4^{x+2}= \ln 5^{?}\]

OpenStudy (anonymous):

\[\ln 4^{x+2}=\ln 5^{x}\]

OpenStudy (anonymous):

\[(x+2)*\ln x = x \ln 5 \]

OpenStudy (anonymous):

(x+2) ln 4 = x ln 5

OpenStudy (anonymous):

(x+2)/x = ln 5 / ln 4

OpenStudy (anonymous):

(x+2) ln 4 = x ln 5 wait, distribute 4 ln x + 2 ln 4 = x ln 5

OpenStudy (anonymous):

i mean x ln 4 + 2 ln 4 = x ln 5 , subtract x ln 4 from both sides

OpenStudy (anonymous):

2 ln 4 = x ln 5 - x ln 4 2 ln 4 = x ( ln 5 - ln 4)

OpenStudy (anonymous):

2ln 4 / ( ln 5 - ln 4 ) = x

OpenStudy (anonymous):

Do it using logs pleeeeeese

OpenStudy (anonymous):

x = 12. 425

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its the same idea,

OpenStudy (anonymous):

just write log instead of ln

OpenStudy (anonymous):

no thankyou , nothing? make me a fan